# Anti-rotors and algebra invariants

### v35

@@ Appendix A Upper triangular Toeplitz matrix algebras
Let $\bf z$} be the $(n\times n)$-matrix with entries having value $1$ on its superdiagonal (the diagonal immediately to the right of the main diagonal) and value 0 otherwise. It is easily shown that for $j<n$ its $j$-th power, ${\bf z}^{j}$, is the $(n\times n)$-matrix with entries $1$ on the $j$-th diagonal to the right of the main diagonal, with all other entries being 0. Thus, we can express the matrix format of $s\in\mathbb{T}_{n}$ as a polynomial in $\bf z$}, such that
@@
We will show that upper triangular Toeplitz matrix algebras have unital norms sensitive to all element components, providing a great contrast with the usual norm for these algebras, which is sensitive to only one element component. One can easily generate other algebras with unital norms “similarly dominating” the usual norm, as seen later.
@@
We must first express $s^{-1}$ in terms of the components of $s=(x_{1},\dots,x_{n})\in\mathbb{T}_{n}$ with $x_{1}\neq 0$. From Proposition A.2, $s^{-1}$ is an upper triangular Toeplitz matrix. But as an $n$-vector, it is easily seen that we can write it as
@@ Appendix B Extension to certain not associative algebras
The two-dimensional real unital algebras (there are three, up to an isomorphism) have generalizations as the Cayley-Dixon sequence algebras ($\mathbb{R}$ and $\mathbb{C}$ being its lowest dimensional members), the Spin Factor Jordan algebras (generalizing $\not{\mathbb{C}}$), and real upper triangular matrix algebras with at least some redundant diagonal components (generalizing the lowest dimensional case, $\mathbb{D}$). Their unital norm groups will be the subject of this subsection.
@@ B.2. The Cayley-Dixon sequence
From the proof of Proposition B.2, it is easily shown that a normalized uncurling metric associated with an algebra of the Cayley-Dixon sequence (all of whose members are -algebras whose Hermitian elements are real) is given by ${\rm diag}\{1,-1,-1,\dots,-1\}$, with the associated special unital norm seen to be the Euclidean norm. Although it is easily shown from results in Section 7 that $\mathbb{C}$ is associated with a 1-parameter family of normalized uncurling metrics, a simple argument (using Lemma B.1) indicates that all other algebras of the sequence have only the above normalized uncurling metric, i.e., the special unital norm family has only one member.
B.3. Spin Factor Jordan Algebras
@@
This makes the vector space $\mathbb{R}\oplus\mathbb{R}^{m}$ a unital algebra with the multiplicative identity $\bf 1$} as simply $1$ (where in this instance $1$ is understood to be an element of the algebra rather than a member of the field). For $m\geq 2$, this is known as a Spin Factor Jordan Algebra [13].

### v36

@@ Appendix A Upper triangular Toeplitz matrix algebras
Let $\bf z$} be the $(n\times n)$-matrix with entries having value $1$ on its superdiagonal (the diagonal immediately to the right of the main diagonal) and value 0 otherwise. It is easily shown that for $j<n$ its $j$-th power, ${\bf z}^{j}$, is the $(n\times n)$-matrix with entries $1$ on the $j$-th diagonal to the right of the main diagonal, with all other entries being 0. Thus, we can express the matrix format of $s\in\mathbb{T}_{n}$ as a polynomial in $\bf z$}, such that,
@@
We will show that upper triangular Toeplitz matrix algebras have unital norms sensitive to all element components, providing a great contrast with the usual norm for these algebras, which is sensitive to only one element component. One can easily generate other algebras with unital norms “similarly dominating” the usual norm, as seen in some other examples in Section 7.
@@
We must first express $s^{-1}$ in terms of the components of $s=(x_{1},\dots,x_{n})\in\mathbb{T}_{n}$ with $x_{1}\neq 0$. From Proposition A.2, $s^{-1}$ is an upper triangular Toeplitz matrix. But as an $n$-vector, it is easily seen from (A.1) that we can write it as
@@ Appendix B Extension to certain not associative algebras
There are three isomorphism classes of two-dimensional real unital algebras, represented by $\mathbb{C}$, $\not{\mathbb{C}}$, and $\mathbb{D}$. The complex numbers $\mathbb{C}$ can be thought of as being generalized by the Cayley-Dixon sequence of algebras. The split-complex numbers $\not{\mathbb{C}}$ can be thought of as being generalized by the formally real Spin Factor Jordan algebras. The dual numbers $\mathbb{D}$ can be thought of as being generalized by the real upper triangular matrix algebras with at least some redundant diagonal components. These generalizations of $\mathbb{D}$ are associative algebras, and have been treated in Appendix A and some examples in Section 7. However, the above generalizations of $\mathbb{C}$ and $\not{\mathbb{C}}$ are not associative (apart from the quaternions). Application of our program to these latter algebras will be presented in this appendix.
@@ B.2. The Cayley-Dixon sequence
From the proof of Proposition B.2, it is easily shown that a normalized uncurling metric associated with an algebra of the Cayley-Dixon sequence (all of whose members are -algebras whose Hermitian elements are real) is given by ${\rm diag}\{1,-1,-1,\dots,-1\}$, with the associated special unital norm seen to be the Euclidean norm. It is easily shown from results in Section 7 that $\mathbb{C}$ is associated with a 1-parameter family of normalized uncurling metrics, which follows from the fact that the space of uncurling metrics is two dimensional. But employing the $\frac{n(n-1)}{2}$ constraint equations with $n>2$, in the context of Lemma B.1, a simple argument indicates that all other algebras of the sequence have only a 1-dimensional anti-rotor. It follows that there is only the single normalized uncurling metric given above.
B.3. Formally real Spin Factor Jordan Algebras
@@
This makes the vector space $\mathbb{R}\oplus\mathbb{R}^{m}$ a unital algebra with the multiplicative identity $\bf 1$} as simply $1$ (where in this instance $1$ is understood to be an element of the algebra rather than a member of the field). For $m\geq 2$, this is known as a (formally real) Spin Factor Jordan Algebra [13].